TESTING THE LIMITS OF INTEGRATION TIME (AND OBSERVER PATIENCE): DETECTION OF IN-TERSTELLAR KCN

\underline{C. SAVAGE}, L.M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721; M. GUÉLIN, Institut de RadioAstronomie Millimétrique, 300 rue de la Piscine, Domaine Universitaire, 38406 Saint Martin d’Hères, France.

KCN has been detected in the circumstellar envelope of IRC+10216, a late-type carbon star, with the Kitt Peak 12meter radio telescope. KCN is a T-shaped, closed shell molecule with a typical asymmetric top spectrum. Five K_{a} = 0 components were detected in the λ = 2-3.5 mm wavelength range. All five lines were fit to a linear rotational diagram, indicating that the KCN emission results from a uniform rotational temperature of 27 K. The derived column density of KCN is ˜ 1 x10^{12} cm^{−2}, corresponding to a fractional abundance of ˜ 3 x10^{−10} relative to H_{2}. All of the lines observed had intensities of 1-2 mK, making them the weakest yet detected by mm-wave astronomy. In order to achieve 3σ detections, integration times for all lines were at least 50 hours, with some frequencies requiring over 100 hours to attain sufficient signal-to-noise. These observations demonstrate the stability of the 12-meter over long periods of time, and confirm the theoretical noise behavior as given by the radiometer equation:

T_{rms} = \frac{2T_{sys}}{η_{spec}\sqrt{Δνt_{scan}}}